# RB-SCD: A New Benchmark for Semantic Change Detection of Roads and Bridges in Traffic Scenes

### v1

@@ Abstract
Accurate detection of changes in roads and bridges, such as construction, renovation, and demolition, is essential for urban planning and traffic management. However, existing methods often struggle to extract fine-grained semantic change information due to the lack of high-quality annotated datasets in traffic scenarios. To address this, we introduce the Road and Bridge Semantic Change Detection (RB-SCD) dataset, a comprehensive benchmark comprising 260 pairs of high-resolution remote sensing images from diverse cities and countries. RB-SCD captures 11 types of semantic changes across varied road and bridge structures, enabling detailed structural and functional analysis. Building on this dataset, we propose a novel framework, Multimodal Frequency-Driven Change Detector (MFDCD), which integrates multimodal features in the frequency domain. MFDCD includes a Dynamic Frequency Coupler (DFC) that fuses hierarchical visual features with wavelet-based frequency components, and a Textual Frequency Filter (TFF) that transforms CLIP-derived textual features into the frequency domain and applies graph-based filtering. Experimental results on RB-SCD and three public benchmarks demonstrate the effectiveness of our approach.
keywords:
[…]
1 Introduction
@@
In the field of remote sensing (RS), change detection (CD) has become an essential task for monitoring environmental, urban, and infrastructure changes (Pekel et al., 2016; Song et al., 2018). With the acceleration of urbanization and the continuous development of transportation infrastructure, CD in traffic scenes is crucial for traffic planning and facility management. In traffic planning, accurately identifying structural and functional changes in roads and bridges helps urban authorities to assess the current state of transportation networks, forecast future demands, and make data-driven decisions about expansions, detours, and upgrades (Lin and Habib, 2022; K S et al., 2024; Wang et al., 2025). For instance, the emergence of new overpasses, flyovers, or road expansions needs to be timely integrated into transportation maps to ensure up-to-date routing and navigation services. Moreover, detecting illegal road constructions or temporary blockages is also critical for enforcing urban regulations and minimizing congestion. From the perspective of traffic facility management, road and bridge CD enables timely identification of structural deterioration, unauthorized demolitions, or unreported construction activities. This semantic-level understanding is crucial for preventive maintenance, disaster response (e.g., after floods or earthquakes), and resource allocation. Additionally, as smart cities and autonomous transportation systems increasingly rely on accurate and up-to-date digital maps, the need for automated and fine-grained CD becomes even more pressing. Roads and bridges are not only physical assets but also key elements in the digital infrastructure of intelligent transportation systems.
@@
However, as shown in Table  1, traditional binary change detection (BCD) datasets (Ji et al., 2018; Chen and Shi, 2020; Shi et al., 2021), etc. primarily focus on building CD and only provide simple information about whether a change has occurred, represented by 0 and 1 for no change and change, respectively. On the other hand, existing semantic change detection (SCD) datasets (Caye Daudt et al., 2019; Yuan et al., 2022; Yang et al., 2022a), etc. mainly focus on changes in land cover, such as building changes and vegetation changes. Changes related to roads and bridges in traffic scenes are either limited in number or completely absent. To address this issue, as annotation examples shown in Figure 1, we propose a new dataset specifically designed for SCD of roads and bridges in traffic scenes: Road and Bridge Semantic Change Detection (RB-SCD) dataset. RB-SCD provides semantic annotations for 11 distinct categories of road and bridge changes, effectively addressing insufficient granularity of information and lack of specificity in detecting road and bridge changes in traffic scenes.
@@
In addition to the RB-SCD dataset, we have also developed a new framework to further facilitate CD performance, the Multimodal Frequency-Driven Change Detector (MFDCD), which integrates multimodal features in the frequency domain. Considering CD methods, traditional approaches often rely on convolutional neural network architectures (Caye Daudt et al., 2018; Fang et al., 2022; Feng et al., 2022), which are insufficient to fully exploit the rich spatiotemporal information in images. Some studies (Chen et al., 2022; Bandara and Patel, 2022; Feng et al., 2023) have incorporated transformers into the network for global modeling. However, these methods still face challenges in accurately capturing and effectively distinguishing complex textures and structural changes. Research has also shown that relying solely on single-modal data has inherent limitations. Recently, some methods  (Yang et al., 2022b; Zhang et al., 2023; Dong et al., 2024) have introduced multimodal features to overcome the shortcomings of single modalities, such as combining visual and textual information or integrating lidar and hyperspectral data. However, these approaches often fail to fully exploit the characteristics of multimodal features, focusing only on modeling the commonalities of multimodal data while neglecting the distributional differences among different modalities.
Table 1: The public datasets for urban change detection with optical remote sensing imagery.
@@
Dataset Resolution (m) Image pairs Image size (pixels) Change type Objects of interest Classes
LMR (Low-to-Medium Resolution)
OSCD  (Daudt et al., 2018) 10 24 $600\times 600$ Binary Land cover
S2MLTCP  (Leenstra et al., 2021) 10 1520 $600\times 600$ Binary Land cover
ZY3  (Zhang and Shi, 2020) 5.8 1 $1154\times 740$ Binary Land cover
Landsat-SCD  (Yuan et al., 2022) 30 8468 $416\times 416$ Semantic Land cover 4
HR (High-Resolution)
WV3  (Zhang and Shi, 2020) 2 2 $1431\times 1431$ Binary Land cover
QB  (Zhang and Shi, 2020) 2.4 1 $1154\times 740$ Binary Land cover
VHR (Very High-Resolution)
AICD  (Bourdis et al., 2011) 0.5 1000 $800\times 600$ Binary Land cover
LEVIR-CD  (Chen and Shi, 2020) 0.5 637 $1024\times 1024$ Binary Buildings
WHU  (Ji et al., 2018) 0.2 1 $32207\times 15354$ Binary Buildings
CDD  (Lebedev et al., 2018) 0.03 - 1 16000 $256\times 256$ Binary Buildings
SYSU  (Shi et al., 2021) 0.5 16000 $256\times 256$ Binary Buildings
BDD  (Adriano et al., 2021) 1.0 1147 $250\times 250$ Semantic Buildings 1
HRSCD  (Caye Daudt et al., 2019) 0.5 291 $10000\times 10000$ Semantic Land cover 5
SECOND  (Yang et al., 2022a) 0.5 - 3 4662 $512\times 512$ Semantic Land cover 6
Hi-UCD  (Tian et al., 2020) 0.1 1293 $1024\times 1024$ Semantic Land cover 9
@@
Figure 1: Annotation examples on RB-SCD dataset. $T_{1}$ and $T_{2}$ represents pre- and post-change images, respectively.
@@
In this work, as shown in Figure 2, we observe that the decomposition and utilization of features in the frequency domain remain underexplored, particularly the fine-grained change information embedded in the high-frequency visual features, the large-scale distribution information in the low-frequency features, and the frequency-domain representation of textual features. The potential patterns of multimodal information in the frequency domain have not been adequately studied. Therefore, we make the first attempt to develop a new architecture that maps image and textual features into different types of frequency domains based on their characteristics, enabling multimodal frequency-domain modeling for CD.
@@
TFF introduces a pre-trained CLIP (Radford et al., 2021) model to obtain textual information, which is further expanded to get text descriptors. Text features contain semantic information but lack spatial features, and discrete Fourier transform (DFT) can convert them into frequency domain information, making it easier to align with the frequency information of image features. By constructing graph filters to aggregate key information, the processed textual features are deeply integrated with visual features to provide multimodal support for SCD.
@@
Our contributions are as follows: {adjustwidth}1.5em0pt•  A comprehensive benchmark for SCD in traffic scenarios. We construct a new dataset named RB-SCD, which focuses on SCD of roads and bridges. It contains 260 large-scale high-resolution RS image pairs, covering over ten cities across multiple countries and annotating 11 typical semantic change types. With its wide geographic diversity and complex traffic infrastructures, RB-SCD supports fine-grained structural and functional change analysis for transportation environments. {adjustwidth}1.5em0pt•  A novel multimodal frequency-driven CD framework. We propose Multimodal Frequency-Driven Change Detector (MFDCD), the first framework to introduce frequency-domain multimodal feature integration for CD. It consists of two main modules: (1) Dynamic Frequency Coupler (DFC), which fuses hierarchical visual features with high- and low-frequency components obtained via wavelet transform to enhance cross-temporal structural change perception; (2) Textual Frequency Filter (TFF), which transforms textual features extracted by a pre-trained CLIP model into the frequency domain using Fourier transform, applies graph-based filtering to extract salient components, and integrates them with visual features for collaborative multimodal representation learning. {adjustwidth}1.5em0pt•  Comprehensive experiments on multiple benchmarks. Extensive experiments on the proposed RB-SCD and three public CD datasets demonstrate that MFDCD achieves superior performance and generalization ability in identifying complex semantic changes in traffic scenes.
2 Related Work
2.1 CD Datasets and Methods
@@
Existing datasets, as shown in Table 1, such as  (Ji et al., 2018; Lebedev et al., 2018; Chen and Shi, 2020; Shi et al., 2021), are primarily designed for BCD tasks and cannot provide specific semantic change information. For SCD, datasets like  (Caye Daudt et al., 2019; Yuan et al., 2022; Yang et al., 2022a) have established some benchmarks. However, these datasets mainly focus on changes in land cover, with little attention paid to the changes in roads and bridges in traffic scenes. In contrast, our RB-SCD dataset is a high-resolution, fine-grained SCD dataset specifically curated and annotated for road and bridge changes in traffic scenes. It captures a wide range of semantic change types and includes diverse geographical contexts, enabling precise and meaningful evaluation of CD algorithms in real-world urban transportation scenarios. By addressing both the lack of semantic detail and the absence of domain specificity in existing datasets, RB-SCD provides a valuable and targeted benchmark for infrastructure monitoring, intelligent transportation analysis, and traffic scene understanding.
@@
Table 2: Specific subcategories in RB-SCD dataset.
@@
In recent years, deep learning has significantly improved the accuracy of CD through its ability to extract relevant features from raw data. Changemask (Zheng et al., 2022) decouples SCD into temporal semantic segmentation and BCD, and combines these two tasks into a unified encoder-transformer-decoder framework. MDENet (Liu et al., 2024) uses MDCM to capture object features in frequency and spatial domains across times, and MFGatt to reduce style impact, focusing on object location. JFRNet (Chang et al., 2024) introduces a three-branch network that takes joint bi-temporal features as unified input, enabling each branch to perceive the semantic relevance of the two time points without any additional interaction operations. DecoderFusion (Li et al., 2024) enhances CD decoding features by leveraging the feature differences between two semantic segmentation decoding branches. CdSC (Wang et al., 2024) introduces a cross-differential semantic consistency network that explores the differences in bi-temporal features while maintaining their semantic consistency. HGINet (Long et al., 2024) builds a multi-level perception aggregation network with a pyramid architecture to extract features that distinguish different categories across multiple levels. However, these methods only consider single-modal data, i.e., images, while neglecting the rich semantic information embedded in multimodal data, thus encountering a performance bottleneck.
2.2 Multimodal Learning
@@
Recent studies have shown that multimodal information is beneficial for visual image analysis. LAVT (Yang et al., 2022b) and VLT (Ding et al., 2023) integrate textual features into visual features through language encoders. Zhang et al. (Zhang et al., 2023) propose a convolutional neural network method with an attention mechanism to enhance feature extraction for lidar data. Their approach leverages the elevation information from lidar data and the spatial-spectral information of hyperspectral data for multimodal feature fusion. MedCLIP (Koleilat et al., 2024) uses text prompts to generate segmentation for clinical scans, adopting a novel decoupled hard-negative noise contrastive estimation loss to fine-tune the BiomedCLIP model. ChangeCLIP (Dong et al., 2024) designs a novel CD framework that utilizes semantic information from image-text pairs and introduce an innovative differential feature compensation module to capture fine-grained semantic changes between them.
2.3 Frequency Transform in Deep Learning
@@
In image processing, the combination of CNN and frequency domain information has been widely applied in tasks such as image restoration, super-resolution, and semantic segmentation. W-Unet (Zhao et al., 2021) replaces down-sampling operations with DWT to reduce the impact of image noise while avoiding information loss. FFT-ReLU  (Mao et al., 2023) performs image deblurring tasks in the frequency domain based on fast Fourier transform. ARFFT (Zhu et al., 2023) proposes a spatial-frequency fusion block to enhance the representation capability of Transformers and extend the receptive field across the entire image. XNet (Zhou et al., 2023) utilizes wavelet transforms to transfer visual information to the frequency domain, leveraging high- and low-frequency information to design a framework supporting both fully supervised and semi-supervised tasks. Given the characteristics of RS images, such as high resolution and susceptibility to spectral variations caused by seasonal changes, using only frequency domain features often results in the loss of critical spatial information. Therefore, in this paper, we propose a novel architecture that models multimodal features in the frequency domain, effectively balancing the advantages of visual features, textual features, and frequency domain features, achieving promising results.
@@
3 RB-SCD Dataset
3.1 Data Annotation and Analysis
Rich Semantic Annotations: As shown in Figure 1 and Figure 3, unlike traditional BCD datasets, we have defined eight major categories (with specific subcategories detailed in Table 2 ), including road, bridge, bare land, building, farmland, water, grass and vegetation. The RB-SCD dataset provides detailed semantic change annotations for 11 distinct types of semantic changes and a background category across these eight categories.
Figure 6: Overview of our proposed Multimodal Frequency-Driven Change Detector (MFDCD).

### v2

@@ Abstract
With the rapid modernization of urban transportation, the accurate detection of changes such as construction, renovation, and demolition of roads and bridges plays a crucial role in urban planning and traffic management. However, existing methods often struggle to precisely extract semantic change information in such scenarios. A major limitation is the scarcity of fine-grained annotated change detection (CD) datasets in traffic scenes, which hinders the training and optimization of relevant algorithms. To address this, we introduce Road and Bridge Semantic Change Detection (RB-SCD) dataset, a comprehensive benchmark containing 260 pairs of high-resolution remote sensing images. The dataset spans a wide range of geographic regions and includes diverse road and bridge types across more than ten cities in multiple countries, incorporating 11 distinct categories of semantic changes, which collectively enable detailed structural and functional change analysis. Leveraging the rich diversity and complexity of RB-SCD, we further propose a novel framework, Multimodal Frequency-Driven Change Detector (MFDCD), which integrates multiple characteristics of multimodal features in the frequency domain for the first time. Technically, MFDCD consists of two components: Dynamic Frequency Coupler (DFC) and the Textual Frequency Filter (TFF). The DFC deeply couples hierarchical visual features with wavelet-derived frequency components, thereby enhancing the perception of detailed and cross-temporal global structural changes. The TFF converts textual features extracted by the pre-trained CLIP model into the frequency domain via Fourier transform. Then it applies graph-based filtering to extract salient frequency components, which are fused with visual features to enable joint multimodal representation learning. Experimental results show that MFDCD performs well on RB-SCD and three public datasets. The RB-SCD dataset, with its rich semantic annotations, serves as a challenging yet valuable resource for road and bridge change analysis in traffic environments. The code and dataset will be available at https://github.com/DaGuangDaGuang/RB-SCD.
Index Terms:
Change detection (CD), road and bridge semantic change detection dataset, dynamic frequency coupler, textual frequency filter, remote sensing (RS), semantic map guidance.
I Introduction
@@
In the field of remote sensing (RS), change detection (CD) has become an essential task for monitoring environmental, urban, and infrastructure changes [1, 2]. With the acceleration of urbanization and the continuous development of transportation infrastructure, CD in traffic scenes is crucial for traffic planning and facility management. In traffic planning, accurately identifying structural and functional changes in roads and bridges helps urban authorities to assess the current state of transportation networks, forecast future demands, and make data-driven decisions about expansions, detours, and upgrades [3][4][5]. For instance, the emergence of new overpasses, flyovers, or road expansions needs to be timely integrated into transportation maps to ensure up-to-date routing and navigation services. Moreover, detecting illegal road constructions or temporary blockages is also critical for enforcing urban regulations and minimizing congestion. From the perspective of traffic facility management, road and bridge CD enables timely identification of structural deterioration, unauthorized demolitions, or unreported construction activities. This semantic-level understanding is crucial for preventive maintenance, disaster response (e.g., after floods or earthquakes), and resource allocation. Additionally, as smart cities and autonomous transportation systems increasingly rely on accurate and up-to-date digital maps, the need for automated and fine-grained CD becomes even more pressing. Roads and bridges are not only physical assets but also key elements in the digital infrastructure of intelligent transportation systems.
@@
However, as shown in Tab. I, traditional binary change detection (BCD) datasets [6][7][8], etc. primarily focus on building CD and only provide simple information about whether a change has occurred, represented by 0 and 1 for no change and change, respectively. On the other hand, existing semantic change detection (SCD) datasets [9][10][11], etc. mainly focus on changes in land cover, such as building changes and vegetation changes. Changes related to roads and bridges in traffic scenes are either limited in number or completely absent. Most existing road extraction datasets [12][13][14] mainly focus on binary pixel segmentation of roads and cannot be used to focus on road changes. To address this issue, as annotation examples shown in Fig. 1, we propose a new dataset specifically designed for SCD of roads and bridges in traffic scenes: Road and Bridge Semantic Change Detection (RB-SCD) dataset. RB-SCD provides semantic annotations for 11 distinct categories of road and bridge changes, effectively addressing insufficient granularity of information and lack of specificity in detecting road and bridge changes in traffic scenes.
@@
In addition to the RB-SCD dataset, we have also developed a new framework to further facilitate CD performance, the Multimodal Frequency-Driven Change Detector (MFDCD), which integrates multimodal features in the frequency domain. Considering CD methods, traditional approaches often rely on convolutional neural network architectures [15][16][17], which are insufficient to fully exploit the rich spatiotemporal information in images. Some studies [18][19][20] have incorporated transformers into the network for global modeling. However, these methods still face challenges in accurately capturing and effectively distinguishing complex textures and structural changes. Research has also shown that relying solely on single-modal data has inherent limitations. Recently, some methods [21][22][23] have introduced multimodal features to overcome the shortcomings of single modalities, such as combining visual and textual information or integrating lidar and hyperspectral data. However, these approaches often fail to fully exploit the characteristics of multimodal features, focusing only on modeling the commonalities of multimodal data while neglecting the distributional differences among different modalities.
TABLE I: Public remote sensing change detection and road extraction datasets. For image counting, the unit of change detection datasets is pairs, while that of road extraction is pieces/images.
@@
Dataset Resolution (m) Image count Image size (pixels) Type Objects of interest Classes
Change detection
OSCD  [24] 10 24 $600\times 600$ Binary Land cover
S2MLTCP  [25] 10 1520 $600\times 600$ Binary Land cover
ZY3  [26] 5.8 1 $1154\times 740$ Binary Land cover
WV3  [26] 2 2 $1431\times 1431$ Binary Land cover
QB  [26] 2.4 1 $1154\times 740$ Binary Land cover
AICD  [27] 0.5 1000 $800\times 600$ Binary Land cover
LEVIR-CD  [7] 0.5 637 $1024\times 1024$ Binary Buildings
WHU  [6] 0.2 1 $32207\times 15354$ Binary Buildings
CDD  [28] 0.03 - 1 16000 $256\times 256$ Binary Buildings
SYSU  [8] 0.5 20000 $256\times 256$ Binary Buildings
BDD  [29] 1.0 1147 $250\times 250$ Semantic Buildings 1
Landsat-SCD  [10] 30 8468 $416\times 416$ Semantic Land cover 4
HRSCD  [9] 0.5 291 $10000\times 10000$ Semantic Land cover 5
SECOND  [11] 0.5 - 3 4662 $512\times 512$ Semantic Land cover 6
Hi-UCD  [30] 0.1 1293 $1024\times 1024$ Semantic Land cover 9
Road extraction
CHN6-CUG  [12] 0.5 4511 $512\times 512$ Binary Road
Massachusetts  [13] 1.5 1000 $1500\times 1500$ Binary Road
Google Earth  [14] 1.2 224 $512\times 512$ Binary Road
@@
Figure 1: Annotation examples on RB-SCD dataset. $T_{1}$ and $T_{2}$ represent the pre- and post-change images, respectively. Label is the ground truth.
@@
In this work, as shown in Fig. 2, we observe that the decomposition and utilization of features in the frequency domain remain underexplored, particularly the fine-grained change information embedded in the high-frequency visual features, the large-scale distribution information in the low-frequency features, and the frequency-domain representation of textual features. The potential patterns of multimodal information in the frequency domain have not been adequately studied. Therefore, we make the first attempt to develop a new architecture that maps image and textual features into different types of frequency domains based on their characteristics, enabling multimodal frequency-domain modeling for CD.
@@
TFF introduces a pre-trained CLIP [31] model to obtain textual information, which is further expanded to get text descriptors. Text features contain semantic information but lack spatial features, and discrete Fourier transform (DFT) can convert them into frequency domain information, making it easier to align with the frequency information of image features. By constructing graph filters to aggregate key information, the processed textual features are deeply integrated with visual features to provide multimodal support for SCD.
Our contributions are as follows:
  • A comprehensive benchmark for SCD in traffic scenarios. We construct a new dataset named RB-SCD, which focuses on SCD of roads and bridges. It contains 260 large-scale high-resolution RS image pairs, covering over ten cities across multiple countries and annotating 11 typical semantic change types. With its wide geographic diversity and complex traffic infrastructures, RB-SCD supports fine-grained structural and functional change analysis for transportation environments.
@@
  • A novel multimodal frequency-driven CD framework. We propose Multimodal Frequency-Driven Change Detector (MFDCD), the first framework to introduce frequency-domain multimodal feature integration for CD. It consists of two main modules: (1) Dynamic Frequency Coupler (DFC), which fuses hierarchical visual features with high- and low-frequency components obtained via wavelet transform to enhance cross-temporal structural change perception; (2) Textual Frequency Filter (TFF), which transforms textual features extracted by a pre-trained CLIP model into the frequency domain using Fourier transform, applies graph-based filtering to extract salient components, and integrates them with visual features for collaborative representation learning.
  • Comprehensive experiments on multiple benchmarks. Extensive experiments on the proposed RB-SCD and three public CD datasets demonstrate that MFDCD achieves superior performance and generalization ability in identifying complex semantic changes in traffic scenes.
II Related Work
II-A CD Datasets and Methods
@@
Existing datasets, as shown in Tab. I, such as  [6][7][8][28], are primarily designed for BCD tasks and cannot provide specific semantic change information. For SCD, datasets like  [9][10][11] have established some benchmarks. However, these datasets mainly focus on changes in land cover, with little attention paid to the changes in roads and bridges in traffic scenes. In contrast, our RB-SCD dataset is a high-resolution, fine-grained SCD dataset specifically curated and annotated for road and bridge changes in traffic scenes. It captures a wide range of semantic change types and includes diverse geographical contexts, enabling precise and meaningful evaluation of CD algorithms in real-world urban transportation scenarios. By addressing both the lack of semantic detail and the absence of domain specificity in existing datasets, RB-SCD provides a valuable and targeted benchmark for infrastructure monitoring, intelligent transportation analysis, and traffic scene understanding.
@@
TABLE II: Specific subcategories in RB-SCD dataset.
@@
In recent years, deep learning has significantly improved the accuracy of CD through its ability to extract relevant features from raw data. Changemask [32] decouples SCD into temporal semantic segmentation and BCD, and combines these two tasks into a unified encoder-transformer-decoder framework. MDENet [33] uses MDCM to capture object features in frequency and spatial domains across times, and MFGatt to reduce style impact, focusing on object location. JFRNet [34] introduces a three-branch network that takes joint bi-temporal features as unified input, enabling each branch to perceive the semantic relevance of the two time points without any additional interaction operations. DecoderFusion [35] enhances CD decoding features by leveraging the feature differences between two semantic segmentation decoding branches. CdSC [36] introduces a cross-differential semantic consistency network that explores the differences in bi-temporal features while maintaining their semantic consistency. HGINet [37] builds a multi-level perception aggregation network with a pyramid architecture to extract features that distinguish different categories across multiple levels. However, these methods only consider single-modal data, i.e., images, while neglecting the rich semantic information embedded in multimodal data, thus encountering a performance bottleneck.
II-B Multimodal Learning
@@
Recent studies have shown that multimodal information is beneficial for visual image analysis. LAVT [21] and VLT [38] integrate textual features into visual features through language encoders. Zhang et al. [22] propose a convolutional neural network method with an attention mechanism to enhance feature extraction for lidar data. Their approach leverages the elevation information from lidar data and the spatial-spectral information of hyperspectral data for multimodal feature fusion. MedCLIP [39] uses text prompts to generate segmentation for clinical scans, adopting a novel decoupled hard-negative noise contrastive estimation loss to fine-tune the BiomedCLIP model. ChangeCLIP [23] designs a novel CD framework that utilizes semantic information from image-text pairs and introduce an innovative differential feature compensation module to capture fine-grained semantic changes between them.
II-C Frequency Transform in Deep Learning
@@
In image processing, the combination of CNN and frequency domain information has been widely applied in tasks such as image restoration, super-resolution, and semantic segmentation. W-Unet [40] replaces down-sampling operations with DWT to reduce the impact of image noise while avoiding information loss. FFT-ReLU  [41] performs image deblurring tasks in the frequency domain based on fast Fourier transform. ARFFT [42] proposes a spatial-frequency fusion block to enhance the representation capability of Transformers and extend the receptive field across the entire image. XNet [43] utilizes wavelet transforms to transfer visual information to the frequency domain, leveraging high- and low-frequency information to design a framework supporting both fully supervised and semi-supervised tasks. Given the characteristics of RS images, such as high resolution and susceptibility to spectral variations caused by seasonal changes, using only frequency domain features often results in the loss of critical spatial information. Therefore, in this paper, we propose a novel architecture that models multimodal features in the frequency domain, effectively balancing the advantages of visual features, textual features, and frequency domain features, achieving promising results.
@@
III RB-SCD Dataset
III-A Data Annotation and Analysis
Rich Semantic Annotations: As shown in Fig. 1 and Fig. 3, unlike traditional BCD datasets, we have defined eight major categories (with specific subcategories detailed in Tab. II), including road, bridge, bare land, building, farmland, water, grass and vegetation. The RB-SCD dataset provides detailed semantic change annotations for 11 distinct types of semantic changes and a background category across these eight categories.
Figure 6: Overview of our proposed Multimodal Frequency-Driven Change Detector (MFDCD). Firstly, the Dynamic Frequency Coupler (DFC) couples hierarchical visual features extracted by ResNet50 with high- and low-frequency components obtained via wavelet transform to enhance cross-temporal structural change perception. Meanwhile, the Textual Frequency Filter (TFF) transforms textual features extracted by a pre-trained CLIP model into the frequency domain using Fourier transform, applies graph-based filtering to extract salient components, and integrates them with visual features for collaborative representation learning. Finally, encoder features are fed into the decoder to obtain the final semantic change detection results.